Displaying Diary Entries for Building: 98 |
Page 1 of 17 (Total Records: 162)
|
Name: | JMR |
Team: | |
Date: | 8/18/2010 |
Entry: | JFB continued in pit F.3331, where a lot of non-articulated human bones now appear everywhere in the northern part and under the large stone, which was taken out today. Before taking it out, we documented the state of excavation in drawing and photo.
E... click here for more... |
|
Name: | JMR |
Team: | |
Date: | 8/19/2010 |
Entry: | LKH continued fill unit 15194. XHB exposed floor U.15341 further (fill above: 15196). It is still continuing towards west, but considerably sloping downwards in that part. He stopped cleaning the floor to make a section through the fill above (15196) in t... click here for more... |
|
Name: | JMR |
Team: | |
Date: | 8/21/2010 |
Entry: | EMM continued working on the clay ball cluster U.15343 in Space 449. Its outlines are visible now, it is a heap of balls in the eastern part of the trench, fading towards west. IF finished room fill U.15335 in Space 310. At the present level, a seemingly ... click here for more... |
|
Name: | JMR |
Team: | |
Date: | 8/22/2010 |
Entry: | Second last day of excavation for this season. BOD continued excavating room fill (15340) in Spaces 450 and 452, typical fill with a lot of building materials and finds. Building 98 is now excavated fairly deep in the central part, not floor yet and the b... click here for more... |
|
Name: | JMR |
Team: | |
Date: | 8/23/2010 |
Entry: | Last day of excavation in Trench 5 for this year.
EMM continued taking out clay ball cluster U.15343 and removed probably is main part, but still some of it is remaining is Space 449, there was no time to take them out and record. We will protect them... click here for more... |
|
Name: | BOD |
Team: | |
Date: | 8/23/2010 |
Entry: | Last excavation day. After a big push downwards in space 450 yesterday (about 40cm down), today was spent extending the same level to space 452. Then – polishing the buttresses' plaster, smartening out the profiles, emptying meep-holes, etc. In sum, movin... click here for more... |
|
Name: | JMR |
Team: | |
Date: | 7/19/2011 |
Entry: | We are moving pleasingly fast in the trench extension of Trench 5. The only limiting factor to the speed of the excavation is the soil sieving, so we are currently working on three sieves.
JHB and RHB finished U.15351 over Space 454. They went down to th... click here for more... |
|
Name: | JMR |
Team: | |
Date: | 7/24/2011 |
Entry: | Excavation in the units with much soil went slowly today as the two sifting workman were missing. DLG continued U.15358, CMB and DKK continued U.15357. HLS continued going down and finding the cut in pit F.3331. It is now rather clear that there must be a... click here for more... |
|
Name: | BOD |
Team: | |
Date: | 7/24/2011 |
Entry: | Entry for the last 2 days, after my computer finally agreed to let me write my diary.
I started excavating U15361 in space 450, continuing on from the end of season 2011. The unit consists of roomfill, with various inclusions. There were a few large ve... click here for more... |
|
Name: | EMM |
Team: | |
Date: | 7/24/2011 |
Entry: | Today I continued working in space 449. Removing some roomfill (U 15362) in the western part of the space, there turned up some rather big lenses of pale yellowish clay, stretching from the South-West corner (F 5057, 3333) to the northern border of the un... click here for more... |
|
|